OpenStudy (anonymous):

a jet skies company charges a flat fee of $24.00 plus $2.25 per hour to rent a jet ski. another company charges a fee of $23.00 plus $2.50 per hour to rent the same jet ski. using a graphing calculator find the number of hours for which the costs are the same round your answer to the nearest whole hour

OpenStudy (anonymous):

you will first have to make an equation for both companies. lets call the first company " SC1 (ski company 1) and the second one SC2 ( ski company 2). we use " x " for number of hours SC1's equation would be: \[24 + 2.25x\] SC2's equation would be: \[23 + 2.50x\] now on your graphing calculator type for \[Y1= 24 + 2.25x\] and \[Y2 = 23 + 2.50x\] Once you've filled these in, go to the graph and adjust your window if needed. Click the 2ND button followed by the "TRACE" button and select option #5 : intersect. It will ask for 1st curve, just click "enter" Then it will ask for 2nd curve, just enter again. It will then ask: Guess? Click enter once more and it will show you the answer for x and y. x being the amount of hours.